OpenStudy (anonymous):

What is the exact volume of a cylinder that has a radius of 9.2 in. and a height of 3.7 in.? [ ] in^3

OpenStudy (anonymous):

@Molly_Qian

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (anonymous):

983.85

OpenStudy (anonymous):

v=πe^2.h

OpenStudy (anonymous):

r*

OpenStudy (studygurl14):

Formula for volume of cylinder: \(\large V=\pi r^2h\) r = radius h = height

OpenStudy (godlovesme):



OpenStudy (godlovesme):

 @horsegirl325

OpenStudy (godlovesme):

@horsegirl325 got it??

OpenStudy (anonymous):

983.34?@Godlovesme

OpenStudy (godlovesme):

yes :)

OpenStudy (anonymous):

thats the answer?